# Data-informativity conditions for structured linear systems with implications for dynamic networks

### v2

@@ 9 Conclusions
$\cdot\begin{bmatrix}w_{i}\\ w_{\!{\scriptscriptstyle\mathcal{D}}^{c}}\\ x_{\!{\scriptscriptstyle\mathcal{T}}}\end{bmatrix}.$
For analyzing the full expression $\Delta\bar{M}_{j}(q,\theta)\kappa^{[j]}(t)$ we need to add the terms $\Delta\bar{M}_{j}^{\xi}(q,\theta)\xi_{\!{\scriptscriptstyle\mathcal{Y}}}^{[j]}(t)$ and $\Delta\bar{M}_{j}^{u}(q,\theta)u^{[j]}(t)$.
Denote $\nu:=\begin{bmatrix}w_{i}\\ w_{\!{\scriptscriptstyle\mathcal{D}}^{c}}\end{bmatrix}$ and $\chi:=\{x_{\!{\scriptscriptstyle\mathcal{T}}}\cup u^{[j]}\}$ Then we can decompose $\nu$ according to $\nu=\nu_{\chi}+\nu_{\perp\chi}$ and using the fact that $\Delta_{j}^{b}(q,\theta)$ is parametrized independently from $\Delta G_{ji}$ and $\tilde{\Delta}_{j}^{a}$, we can rewrite $\Delta\bar{M}_{j}^{w}(q,\theta)w_{\!{\scriptscriptstyle\mathcal{D}}}^{[j]}$ as
@@
$\cdot\begin{bmatrix}\nu_{\chi}+\nu_{\perp\chi}\\ x_{\!{\scriptscriptstyle\mathcal{T}}}\end{bmatrix}$
The expression for $\Delta\bar{M}_{j}^{u}(q,\theta)w_{\!{\scriptscriptstyle\mathcal{D}}}^{[j]}$ is now written as a summation of two terms that are mutually uncorrelated: one driven by $x_{\!{\scriptscriptstyle\mathcal{T}}}$ and $\nu_{\chi}$ and one driven by a term that it orthogonal to this. For evaluating $\Delta\bar{M}_{j}(q,\theta)\kappa^{[j]}(t)$ two additional terms are added, given by $\Delta\bar{M}_{j}^{\xi}(q,\theta)\xi_{\!{\scriptscriptstyle\mathcal{Y}}}^{[j]}$ and $\Delta\bar{M}_{j}^{u}(q,\theta)u^{[j]}$.
@@
as specified in (33). This implies $\Delta\bar{G}_{ji}(q,\theta)=0$ in the minimum of the quadratic cost function. $\Box$
@@
When applying the graph-based results of Proposition 5 to the results of Theorem theox, the graph-based condition becomes to having $dim(\eta^{[j]})$ vertex disjoint paths from all external signals $(u,e)$ to $\eta^{[j]}$. With the expression for $\eta^{[j]}$ (33) this becomes: $dim(\eta^{[j]})$ vertex disjoint paths from $(u,e)\backslash\chi$ to $(w_{i\cup\!{\scriptscriptstyle\mathcal{D}}^{c}},\xi_{\!{\scriptscriptstyle\mathcal{Y}}}^{[j]})$.
@@
In the vertex disjoint path condition this set of signals $e^{\!{\scriptscriptstyle\mathcal{Y}}}$ is then required to establish the positive definite spectrum condition for $\xi_{\!{\scriptscriptstyle\mathcal{Y}}}^{[j]}$, which is satisfied by definition. The remaining term then becomes: $dim(\!{\scriptscriptstyle\mathcal{D}}^{c})+1$ vertex disjoint paths between $(u,e\backslash e^{\!{\scriptscriptstyle\mathcal{Y}}})\backslash\chi$ and $w_{i\cup\!{\scriptscriptstyle\mathcal{D}}^{c}}$. Or equivalently $dim(\!{\scriptscriptstyle\mathcal{D}}^{c})+1$ vertex disjoint paths between $(u\backslash u^{[j]},e\backslash e^{\!{\scriptscriptstyle\mathcal{Y}}})\backslash x_{\!{\scriptscriptstyle\mathcal{T}}^{*}}$ and $w_{i\cup\!{\scriptscriptstyle\mathcal{D}}^{c}}$, which proves the result.

### v3

@@ 10 Conclusions
$\cdot\begin{bmatrix}w_{i}\\ w_{\!{\scriptscriptstyle\mathcal{D}}^{c}}\\ {\color[rgb]{0,0,0}x_{\!{\scriptscriptstyle\mathcal{T}}^{*}}}\end{bmatrix}$.
Denote $\nu:=\begin{bmatrix}w_{i}\\ w_{\!{\scriptscriptstyle\mathcal{D}}^{c}}\end{bmatrix}$ and $\chi:=\{{\color[rgb]{0,0,0}x_{\!{\scriptscriptstyle\mathcal{T}}^{*}}}\cup u^{[j]}\}$ . Then we can decompose $\nu$ according to $\nu=\nu_{\chi}+\nu_{\perp\chi}$ and using the fact that $\Delta_{j}^{b}(q,\theta)$ is parametrized independently from $\Delta G_{ji}$ and $\tilde{\Delta}_{j}^{a}$, we can rewrite $\Delta\bar{M}_{j}^{w}(q,\theta)w_{\!{\scriptscriptstyle\mathcal{D}}}^{[j]}$ as
@@
$\cdot\begin{bmatrix}\nu_{\chi}+\nu_{\perp\chi}\\ {\color[rgb]{0,0,0}x_{\!{\scriptscriptstyle\mathcal{T}}^{*}}}\end{bmatrix}$
The expression for $\Delta\bar{M}_{j}^{u}(q,\theta)w_{\!{\scriptscriptstyle\mathcal{D}}}^{[j]}$ is now written as a summation of two terms that are mutually uncorrelated: one driven by $\color[rgb]{0,0,0}x_{\!{\scriptscriptstyle\mathcal{T}}^{*}}$} and $\nu_{\chi}$ and one driven by a term that is orthogonal to this. For evaluating $\Delta\bar{M}_{j}(q,\theta)\kappa^{[j]}(t)$ two additional terms are added, given by $\Delta\bar{M}_{j}^{\xi}(q,\theta)\xi_{\!{\scriptscriptstyle\mathcal{Y}}}^{[j]}$ and $\Delta\bar{M}_{j}^{u}(q,\theta)u^{[j]}$.
@@
as specified in (28). This implies that $\Delta\bar{G}_{ji}(q,\theta)=0$ in the minimum of the quadratic cost function. $\Box$
@@
[…]
@@
[…]
Appendix E Proof of Corollary 10.
External excitation signals that are directly entering input nodes $w_{\!{\scriptscriptstyle\mathcal{D}}}$ will contribute to $u^{\perp j}$, except when there is a confounding variable between the concerned input and $w_{\!{\scriptscriptstyle\mathcal{Y}}}$. The input node is then classified to belong to set $w_{\!{\scriptscriptstyle\mathcal{B}}}$, VandenHof&etal_IFAC:23 . For the data-informativity conditions, one can always construct a disconnecting set $w_{\!{\scriptscriptstyle\mathcal{D}}^{c}}$ consisting of nodes that are out-neighbors of $w_{i}$. For satisfying the structural conditions a. and b., nodes in $w_{\!{\scriptscriptstyle\mathcal{B}}}$ are not allowed to be out-neighbors of $w_{i}$. As a result there always exist a disconnecting set that does not contain nodes in $w_{\!{\scriptscriptstyle\mathcal{B}}}$, and external excitation on these nodes will always appear in $u^{\perp j}$.